# HelixFold-Multimer: Elevating Protein Complex Structure Prediction to New Heights

### v1

@@ Abstract
While monomer protein structure prediction tools boast impressive accuracy, the prediction of protein complex structures remains a daunting challenge in the field. This challenge is particularly pronounced in scenarios involving complexes with protein chains from different species, such as antigen-antibody interactions, where accuracy often falls short. Limited by the accuracy of complex prediction, tasks based on precise protein-protein interaction analysis also face obstacles. In this report, we highlight the ongoing advancements of our protein complex structure prediction model, HelixFold-Multimer, underscoring its enhanced performance. HelixFold-Multimer provides precise predictions for diverse protein complex structures, especially in therapeutic protein interactions. Notably, HelixFold-Multimer achieves remarkable success in antigen-antibody and peptide-protein structure prediction, surpassing AlphaFold-Multimer by several folds. HelixFold-Multimer is now available for public use on the PaddleHelix platform, offering both a general version and an antigen-antibody version. Researchers can conveniently access and utilize this service for their development needs.
@@ 1 Introduction
We introduce HelixFold-Multimer, a novel approach that significantly enhances the accuracy and widens the application scope of protein complex structure prediction. Building upon the groundwork laid by our prior work, HelixFold [16] and HelixFold-Single [1], HelixFold-Multimer improves cross-chain interaction modeling by integrating domain expertise into model architecture, input features, and training strategies. HelixFold-Multimer is designed to provide highly accurate predictions for a diverse array of protein complex structures, particularly those involved in therapeutic protein interactions. We find that HelixFold-Multimer exhibits impressive effectiveness in accurately predicting a wide range of complex protein structures, notably excelling in predicting antigen-antibody, nanobody-antigen interfaces, and peptide-protein interfaces. Furthermore, HelixFold-Multimer demonstrates exceptional performance with commonly studied protein targets in drug development, suggesting its potential to optimize the protein design process for therapeutic development. HelixFold-Multimer currently consists of two versions: a general version designed for predicting common protein complex structures, particularly interfaces between peptides and proteins, and a specialized version developed for predicting antigen-antibody structures. Both the general version and the antigen-antibody version are publicly available for use on the PaddleHelix platform.
@@ 2.1.1 Overall Performance of General Proteins
Figure 1: Comparison between AlphaFold and HelixFold-Multimer for heteromeric protein complexes. The box plots on the left show DockQ score distributions, while the bar graph on the right reflects the percentage of accurate (DockQ >>> 0.23).
Figure 2: Comparison of AlphaFold and HelixFold-Multimer in protein-peptide docking accuracy. The box plots on the left show DockQ score distributions, while the bar graph on the right reflects the percentage of accurate (DockQ >>> 0.23).
Heteromeric complexes are formed by the interaction of multiple distinct protein chains. The comparison of HelixFold-Multimer with AlphaFold and RosettaFold on heteromeric protein complexes is summarized in Figure 1. The median DockQ score of HelixFold-Multimer (0.304) is comparable to that of AlphaFold (0.316). When considering percentage of the correct samples with DockQ > 0.23, the performance of HelixFold-Multimer (57.8%) exceeds that of AlphaFold (53.6%) by 4.2%. This highlights the superior capabilities of HelixFold-Multimer in the task of predicting the structures of heteromeric protein complexes.
Peptides, being shorter chains of amino acids compared to proteins, can display a wide range of flexible structures. The challenge in predicting the precise structures of protein-peptide complexes lies in accurately modeling the flexibility of the peptides. The results of protein-peptide interfaces is shown in Figure 2. HelixFold-Multimer achieved a median DockQ score of 0.295, marking a pronounced improvement compared to the median DockQ score of 0.262 obtained by AlphaFold and 0.093 by RoseTTAFold. This enhancement in predictive accuracy is further evidenced by the success rate metric (DockQ > 0.23), where HelixFold-Multimer attained a success rate of 68.9%. In contrast, AlphaFold exhibited a success rate of 54.1% for the same threshold. The structural analysis of protein-peptide complexes profoundly influences the research and development of peptide therapeutics. Understanding the intricate interplay between proteins and peptides not only aids in the design of more effective drugs but also sheds light on fundamental biological processes, paving the way for innovative therapeutic strategies.
@@ 2.2 Results of Antigen-Antibody Version
Antigen-antibody structure prediction is essential for unraveling immune responses and informing drug design, vaccine development, and diagnostic assays. It serves as a cornerstone in advancing biomedical research and enhancing therapeutic interventions. When assessing HelixFold-Multimer for antigen-antibody, we evaluate its effectiveness across three key dimensions: overall performance, model confidence, and efficacy on various antigen-antibody categories.
@@ 2.2.1 Overall Comparison
Figure 3: Overall comparison for antibody-related complex structure predictions. Panels (a), and (b) show the distribution of DockQ scores, which indicate the quality of the predicted structures, for the antibody-antigen interface and nanobody-antigen interface, respectively. Meanwhile, panels (c), and (d), display the percentage of correct predictions for these interfaces. The data underscores a notable improvement in prediction accuracy with the HelixFold-Multimer model compared to both AlphaFold and RoseTTAFold across all interfaces evaluated.
We first assess HelixFold-Multimer on antibody-related complex structure predictions across two categories: Antibody-Antigen Interfaces (consisting of heavy and light chains) and Nanobody-Antigen Interfaces, as illustrated in Figure 3. HelixFold-Multimer exhibits significant advantages in predicting both antibody-antigen and nanobody-antigen interfaces. In the evaluation of antibody-antigen interfaces, HelixFold-Multimer’s predictions surpassed those of AlphaFold, achieving a mean DockQ score of 0.390, marking a substantial 5-fold improvement. HelixFold-Multimer achieves an impressive success rate (DockQ > 0.23) of 52.7% in accurately predicting antibody-antigen structures. This contrasts sharply with AlphaFold’s success rate of 7.6% and RoseTTAFold’s 4.6%. The HelixFold-Multimer outperformed both AlphaFold and RoseTTAFold in predicting the nanobody-antigen interface. It achieved a median DockQ score of 0.703 and a mean DockQ of 0.538, representing a significant improvement over previous baselines. With a success rate of 69.2%, HelixFold-Multimer demonstrated superior predictive capabilities compared to AlphaFold (15.4%) and RoseTTAFold (7.7%). This indicates its effectiveness in accurately predicting interactions, even for smaller and more variable nanobody structures.
Figure 4: Evaluation of antibody VH-VL interfaces: Box plots on the left illustrate the performance comparison among RoseTTAFold, AlphaFold, and HelixFold-Multimer. The bar graph on the right quantifies the percentage of predictions with very high accuracy (DockQ > 0.8).
We subsequently evaluated the predictive accuracy for antibody VH-VL interfaces, as shown in Figure 4. While existing protein structure prediction tools have achieved decent accuracy in predicting the conformation of antibodies (including heavy chain H and light chain L), higher-precision prediction results remain highly meaningful for antibody understanding. The prediction of the VH-VL chain interface by HelixFold-Multimer yielded a median DockQ score of 0.823, surpassing AlphaFold’s score of 0.774 and RoseTTAFold’s score of 0.653. HelixFold-Multimer achieved a very high accuracy rate (DockQ > 0.8) of 59.5%, representing a significant improvement over AlphaFold’s 37.4% and RoseTTAFold’s 19.1%. These results indicate the proficiency of HelixFold-Multimer in accurately modeling antibody structures.
@@ 2.2.2 Model Confidence
The correlation between confidence scores and model accuracy was exhaustively examined to evaluate the reliability of the HelixFold-Multimer in predicting antigen-antibody interactions.
We performed an analysis of three scoring metrics within the HelixFold-Multimer model, which assesses the confidence levels of the inference results: confidence scores (CFS), iPTM scores, and pLDDT scores. Subsequently, we organized the test samples into groups based on these confidence scores and presented the distribution of DockQ scores within each group (see Figure 5(a)-(c)). The figures suggest a clear pattern: in general, antigen-antibody complexes identified by the model as having high confidence levels tend to have significantly higher DockQ scores compared to those with lower confidence levels. Figure 5(d)-(f) elucidates the relationship between the accuracy of test samples and scoring metrics through scatter plots. All the scoring metrics show correlations with DockQ. Notably, confidence scores (Pearson correlation: 0.664) and iPTM (Pearson correlation: 0.658) exhibit stronger correlations with DockQ compared to pLDDT (Pearson correlation: 0.344). This robust correlation between scoring metrics and DockQ scores will serve as valuable guidance for leveraging HelixFold-Multimer in antibody development. For instance, analyzing these scoring metrics can help identify antibodies or antigens with higher research potential.
@@ 2.2.3 Efficacy on Various Antigen-Antibody Categories
We conducted an assessment of HelixFold-Multimer’s effectiveness across various species groups and sequence identity intervals to identify its strengths in specific Antigen-Antibody categories. Species were classified into three groups: Homo sapiens, Mus musculus, and other species (including macaca mulatta, rattus norvegicus, synthetic construct, gallus gallus, and oryctolagus cuniculus), based on sample counts. The classification of species was based on the origin of the antibody heavy chains. The sequence identity was calculated through MMseqs2 [17] on antigen chains.
Different species and sequence similarities exhibit significant differences. As shown in the left subfigure of Figure 6, HelixFold-Multimer achieves higher mean DockQ scores for Homo sapiens (0.43) and Mus musculus (0.43) compared to other species (0.15). This outcome aligns with expectations, given that the majority of antibody research focuses on Homo sapiens and Mus musculus. When evaluating antigens grouped by sequence similarity (the right subfigure of Figure 6, we observe a strong correlation between antigen sequence similarity and accuracy, as expected. HelixFold-Multimer achieves a mean DockQ score of up to 0.56 for samples corresponding to antigens already present in the training dataset. We believe that HelixFold-Multimer’s performance on popular antigen-antibody complexes is already quite reliable and can effectively assist researchers in analyzing interactions between antigens and antibodies. In the future, we aim to further enhance the model’s performance on less common antigens.
@@ 3 Conclusion
HelixFold-Multimer showcases exceptional prowess in predicting the structures of diverse protein complexes, surpassing prior methodologies by a significant margin. Particularly noteworthy is its outstanding accuracy in forecasting the structures of protein-peptide and antigen-antibody complexes, crucial components in therapeutic interventions. These findings not only highlight the model’s potential in advancing drug design for macromolecular therapeutics but also underscore its pivotal role in shaping the future landscape of therapeutic development. By harnessing the predictive power of HelixFold-Multimer, researchers are poised to unlock new possibilities in precision medicine and therapeutic innovation.
@@ 1.1 Baselines
RosettaFold [18] integrates deep learning with biochemical knowledge, showing promising results in single-chain protein structure prediction. However, its accuracy in predicting multi-chain protein complex structures is not satisfactory. On the other hand, AlphaFold [2, 8], particularly its Multimer version [8], AlphaFold-Multimer, performs exceptionally well in predicting general protein complex structures. Nonetheless, its accuracy drops significantly in scenarios lacking sufficient cross-chain evolutionary information, such as antigen-antibody complexes.
For the sake of convenience in comparison, this report contrasts the results of single-model inference rather than using ensemble methods (the original AlphaFold employs inference from 25 models followed by ensemble). We conducted inference using version 2.3 model 1 from AlphaFold’s GitHub repository. For the RoseTTAFold model, we referred to the setup described in RoseTTAFold2 [19].
@@ 1.2 Evaluation Metrics
To evaluate the accuracy of predicted protein complexes, we employed the DockQ metric [20]. Consistent with common practice, we utilized the DockQ threshold to gauge the success rates of model predictions. Predictions with a DockQ value above 0.23 were deemed accurate, while those surpassing 0.8 were categorized as exhibiting very high accuracy.
@@ 1.3 Confidence Metrics
In our study, three scoring metrics were used: pLDDT (the predicted local distance difference test) [21], PTM (the predicted TM-score) [22], and confidence score. pLDDT and PTM are metrics provided by the model and are indicators of overall structural accuracy. For interface accuracy of different chains, we measured via iPTM (the interface predicted TM-score), which is calculated based on the residues at the interface. Additionally, we used the confidence score[8] index intended to simultaneously consider the accuracy of the inter-chain interface and the overall prediction accuracy. The confidence score was calculated by assigning different weights to the PTM and iPTM values.

### v2

@@ Abstract
While monomer protein structure prediction tools boast impressive accuracy, the prediction of protein complex structures remains a daunting challenge in the field. This challenge is particularly pronounced in scenarios involving complexes with protein chains from different species, such as antigen-antibody interactions, where accuracy often falls short. Limited by the accuracy of complex prediction, tasks based on precise protein-protein interaction analysis also face obstacles. In this report, we highlight the ongoing advancements of our protein complex structure prediction model, HelixFold-Multimer, underscoring its enhanced performance. HelixFold-Multimer provides precise predictions for diverse protein complex structures, especially in therapeutic protein interactions. Notably, HelixFold-Multimer achieves remarkable success in antigen-antibody and peptide-protein structure prediction, greatly surpassing AlphaFold 3. HelixFold-Multimer is now available for public use on the PaddleHelix platform, offering both a general version and an antigen-antibody version. Researchers can conveniently access and utilize this service for their development needs.
@@ 1 Introduction
We introduce HelixFold-Multimer, a novel approach that significantly enhances the accuracy and widens the application scope of protein complex structure prediction. Building upon the groundwork laid by our prior work, HelixFold [16] and HelixFold-Single [1], HelixFold-Multimer improves cross-chain interaction modeling by integrating domain expertise into model architecture, input features, and training strategies. HelixFold-Multimer is designed to provide highly accurate predictions for a diverse array of protein complex structures, particularly those involved in therapeutic protein interactions. We find that HelixFold-Multimer exhibits impressive effectiveness in accurately predicting a wide range of complex protein structures, notably excelling in predicting antigen-antibody and peptide-protein interfaces. Delineating the epitope regions of the antigen allows for further improvement in the prediction accuracy of antigen-antibody structure. Furthermore, HelixFold-Multimer demonstrates exceptional performance with commonly studied protein targets in drug development, suggesting its potential to optimize the protein design process for therapeutic development. HelixFold-Multimer currently consists of two versions: a general version designed for predicting common protein complex structures, particularly interfaces between peptides and proteins, and a specialized version developed for predicting antigen-antibody structures. Both the general version and the antigen-antibody version are publicly available for use on the PaddleHelix platform.
@@ 2.1.1 Overall Performance of General Proteins
Figure 1: Comparison between AlphaFold 2.3, AlphaFold 3 and HelixFold-Multimer for heteromeric protein complexes. The box plots on the left show DockQ score distributions, while the bar graph on the right reflects the percentage of accuracy (DockQ >>> 0.23).
Figure 2: Comparison of AlphaFold 2.3, AlphaFold 3 and HelixFold-Multimer in protein-peptide docking accuracy. The box plots on the left show DockQ score distributions, while the bar graph on the right reflects the percentage of accuracy (DockQ >>> 0.23).
Heteromeric complexes are formed by the interaction of multiple distinct protein chains. The comparison of HelixFold-Multimer with AlphaFold 2.3 [8], AlphaFold 3 [17] and RosettaFold [18] on heteromeric protein complexes is summarized in Figure 1. The mean DockQ score of HelixFold-Multimer (0.378) is higher than that of AlphaFold 2.3 (0.316), but lower than AlphaFold 2.3 (0.438). When considering the percentage of the correct samples with DockQ > 0.23, AlphaFold 3 shows a leading performance with 64.9% of its predictions above 0.23, demonstrating its robustness in protein structure prediction compared to AlphaFold 2.3 and HelixFold-Multimer. HelixFold-Multimer, while slightly behind AlphaFold 3, still outperforms AlphaFold 2.3, This result indicates that HelixFold Multimer offers predictive performance on par with AlphaFold in modeling heteromeric protein complexes.
Peptides, being shorter chains of amino acids compared to proteins, can display a wide range of flexible structures. The challenge in predicting the precise structures of protein-peptide complexes lies in accurately modeling the flexibility of the peptides. The results of protein-peptide interfaces is shown in Figure 2. HelixFold-Multimer achieved a median DockQ score of 0.295, marking a pronounced improvement compared to the median DockQ score of 0.260 obtained by AlphaFold 3 and 0.093 by RoseTTAFold. This enhancement in predictive accuracy is further evidenced by the success rate metric (DockQ > 0.23), where HelixFold-Multimer attained a success rate of 68.9%. In contrast, AlphaFold 3 exhibited a success rate of 59.0% for the same threshold. The structural analysis of protein-peptide complexes profoundly influences the research and development of peptide therapeutics. Understanding the intricate interplay between proteins and peptides not only aids in the design of more effective drugs but also sheds light on fundamental biological processes, paving the way for innovative therapeutic strategies.
@@ 2.2 Results of Antigen-Antibody Version
Antigen-antibody structure prediction holds paramount importance in comprehending immune responses and driving advancements in drug design, vaccine development, and diagnostic assays. It stands as a fundamental pillar in the progression of biomedical research and the enhancement of therapeutic interventions.
In evaluating HelixFold-Multimer for its efficacy in predicting antigen-antibody interfaces, we embark on a comprehensive assessment. Initially, we gauge the model’s performance not only on the antigen-antibody interface but also on the interface between the antibody heavy and light chains. Subsequent to this assessment, we undertake a comparative analysis to discern the impact of accurately delineating epitope regions within the specified antigen. Further analysis is undertaken on the model’s confidence and efficacy across different antigen-antibody categories. Finally, we provide a case study analysis.
@@ 2.2.1 Overall Comparison
Figure 3: Overall comparison for antibody-related complex structure predictions. The left figure depicts the distribution of DockQ scores, serving as a metric for the quality of predicted structures. The right figure showcases the percentage of accurate predictions specifically concerning antibody-antigen interfaces. The data underscores a notable improvement in prediction accuracy with the HelixFold-Multimer model compared to RoseTTAFold, AlphaFold 2.3, and AlphaFold 3.
Figure 3 illustrates the results of HelixFold-Multimer and several strong baseline methods on antibody-antigen interfaces. The median DockQ score of HelixFold-Multimer stands at 0.38, which notably outperforms AlphaFold 3 (0.07). Additionally, HelixFold-Multimer demonstrates a notable success rate of 58.8% in accurately predicting antibody-antigen structures. This success rate surpasses that of AlphaFold 3 (38.9%) and RoseTTAFold (5.3%). These results indicate the superior predictive capability of HelixFold-Multimer, offering promising implications for antibody-antigen interaction studies.
Figure 4: Evaluation of antibody VH-VL interfaces: Box plots on the left illustrate the performance comparison among RoseTTAFold, AlphaFold 2.3, AlphaFold 3 and HelixFold-Multimer. The bar graph on the right quantifies the percentage of predictions with very high accuracy (DockQ > 0.8).
We subsequently evaluate the predictive accuracy for antibody VH-VL interfaces, as shown in Figure 4. While existing protein structure prediction tools have achieved decent accuracy in predicting the conformation of antibodies (including heavy chain H and light chain L), higher-precision prediction results remain highly meaningful for antibody understanding. The prediction of the VH-VL chain interface by HelixFold-Multimer and AlphaFold 3 yielded the same median DockQ score of 0.823, surpassing AlphaFold 2.3’s score of 0.774 and RoseTTAFold’s score of 0.653. HelixFold-Multimer achieves a very high accuracy rate (DockQ > 0.8) of 59.5%, slightly higher than Alphafold 3’s 58.5%, representing a significant improvement over AlphaFold 2.3’s 37.4% and RoseTTAFold’s 19.1%. These results indicate the proficiency of HelixFold-Multimer in accurately modeling antibody structures.
@@
2.2.2 Impact of Epitope Specification for Antigen-Antibody Structure Prediction
The designation of antigen epitopes plays a crucial role in predicting antigen-antibody structures. It assists in accurately locating the binding interface, enhancing model accuracy and prediction specificity. Typically, researchers can determine antigen binding sites through analyses such as deep mutational scanning (DMS). We investigate whether there will be further improvements in accuracy after specifying the antigen epitope for HelixFold-Multimer. Figure 5 indicates that after specifying the antigen epitope, there is a further improvement in the prediction accuracy of HelixFold-Multimer.
Figure 5: Impact of epitope specification for antigen-antibody structure prediction.
@@ 2.2.3 Model Confidence
The correlation between confidence scores and model accuracy is exhaustively examined to evaluate the reliability of the HelixFold-Multimer in predicting antigen-antibody interactions.
We perform an analysis of three scoring metrics within the HelixFold-Multimer model, which assesses the confidence levels of the inference results: confidence scores (CFS), iPTM scores, and pLDDT scores. Subsequently, we organize the test samples into groups based on these confidence scores and present the distribution of DockQ scores within each group (see Figure 6(a)-(c)). The figures suggest a clear pattern: in general, antigen-antibody complexes identified by the model as having high confidence levels tend to have significantly higher DockQ scores compared to those with lower confidence levels. Figure 6(d)-(f) elucidates the relationship between the accuracy of test samples and scoring metrics through scatter plots. All the scoring metrics show correlations with DockQ. Notably, confidence scores (Pearson correlation: 0.664) and iPTM (Pearson correlation: 0.658) exhibit stronger correlations with DockQ compared to pLDDT (Pearson correlation: 0.344). This robust correlation between scoring metrics and DockQ scores will serve as valuable guidance for leveraging HelixFold-Multimer in antibody development. For instance, analyzing these scoring metrics can help identify antibodies or antigens with higher research potential.
@@ 2.2.4 Efficacy on Various Antigen-Antibody Categories
We conduct an assessment of HelixFold-Multimer’s effectiveness across various species groups and sequence identity intervals to identify its strengths in specific Antigen-Antibody categories. Species were classified into three groups: Homo sapiens, Mus musculus, and other species (including macaca mulatta, rattus norvegicus, synthetic construct, gallus gallus, and oryctolagus cuniculus), based on sample counts. The classification of species is based on the origin of the antibody heavy chains. The sequence identity is calculated by MMseqs2 [19] on antigen chains.
Different species and sequence similarities exhibit significant differences. As shown in the left subfigure of Figure 7, HelixFold-Multimer achieves higher mean DockQ score for Homo sapiens (0.43) and Mus musculus (0.43) compared to other species (0.15). This outcome aligns with expectations, given that the majority of antibody data originate from the research of Homo sapiens and Mus musculus. When evaluating antigens grouped by sequence similarity (the right subfigure of Figure 7, we observe a strong correlation between antigen sequence similarity and accuracy, as expected. HelixFold-Multimer achieves a mean DockQ score of up to 0.56 for samples corresponding to antigens already present in the training dataset. We believe that HelixFold-Multimer’s performance on popular antigen-antibody complexes is already quite reliable and can effectively assist researchers in analyzing interactions between antigens and antibodies. In the future, we aim to further enhance the model’s performance on less popular targets.
2.2.5 Case Study
We conduct visualizations using two significant viral pathogens, namely HIV and SARS-CoV-2, as benchmarks to compare the predictive capabilities of HelixFold-Multimer and AlphaFold 3. In both instances, HelixFold-Multimer exhibits exceptional performance by accurately pinpointing the binding interfaces and offering highly precise structural predictions. This precision underscores its ability to precisely capture the complex interplay between the viral antigens and antibody proteins. Conversely, AlphaFold 3’s predictions fall short, as it either failed to identify the binding interfaces altogether or produced structural deviations, indicating potential limitations in its ability to accurately model antigen-antibody interactions.
Figure 8: Comparison of Antibody-Antigen Predictions by HelixFold-Multimer and AlphaFold 3. This figure illustrates the structural predictions for two antibody-antigen complexes: one with HIV (8SAQ) and the other with SARS-CoV-2 (7XS8). Each set shows a comparison between the actual structure (ground truth). HelixFold-Multimer demonstrating closer alignment to the true structures compared to AlphaFold 3.
@@ 3 Conclusion
HelixFold-Multimer showcases exceptional prowess in predicting the structures of diverse protein complexes, surpassing prior methodologies by a significant margin. Particularly noteworthy is its outstanding accuracy in forecasting the structures of protein-peptide and antigen-antibody complexes, crucial components in therapeutic interventions. These findings not only highlight the model’s potential in advancing drug design for macromolecular therapeutics but also underscore its pivotal role in shaping the future landscape of therapeutic development. By harnessing the predictive power of HelixFold-Multimer, researchers are poised to unlock new possibilities in therapeutic innovation.
@@ 1.1 Baselines
RosettaFold [20] integrates deep learning with biochemical knowledge, showing promising results in single-chain protein structure prediction. However, its accuracy in predicting multi-chain protein complex structures is not satisfactory. On the other hand, AlphaFold [2, 8], particularly its Multimer version [8], AlphaFold-Multimer, performs exceptionally well in predicting general protein complex structures. Nonetheless, its accuracy drops significantly in scenarios lacking sufficient cross-chain evolutionary information, such as antigen-antibody complexes. The recent released Alphafold 3[17] largely improved the predictive performance of antigen-antibody complex.
For the sake of convenience in comparison, this report contrasts the results of single-model inference rather than using ensemble methods (the original AlphaFold 2.3 employs inference from 25 models followed by ensemble). We conducte inference using version 2.3 model 1 from AlphaFold’s GitHub repository. For AlphaFold 3, we use the official online server to generate predictions and calculate metrics based on the model 0 result that has the highest confidence score among the five predicted outcomes. For the RoseTTAFold model, we refer to the setup described in RoseTTAFold2 [21].
@@ 1.2 Evaluation Metrics
To evaluate the accuracy of predicted protein complexes, we employ the DockQ metric [22]. Consistent with common practice, we utilize the DockQ threshold to gauge the success rates of model predictions. Predictions with a DockQ value above 0.23 were deemed accurate, while those surpassing 0.8 were categorized as exhibiting very high accuracy.
@@ 1.3 Confidence Metrics
In our study, three scoring metrics are used: pLDDT (the predicted local distance difference test) [23], PTM (the predicted TM-score) [24], and confidence score. pLDDT and PTM are metrics provided by the model and are indicators of overall structural accuracy. For interface accuracy of different chains, we measure via iPTM (the interface predicted TM-score), which is calculated based on the residues at the interface. Additionally, we use the confidence score[8] index intended to simultaneously consider the accuracy of the inter-chain interface and the overall prediction accuracy. The confidence score is calculated by assigning different weights to the PTM and iPTM values.